The relative frequency distribution of the number of phobias reported by a hypothetical sample of 500 college students is given as follows. 0–2 0.47 3–5 0.25 6–8 0.14 9–11 0.09 12–14 0.05 (a) What is the probability that a college student expresses fewer than three phobias? (b) What is the probability that a college student has between 3 and 11 phobias?
Answer #1

a) P(Fewer than three) = P(0-2) = 0.47

b) P(Between 3 and 11) = 0.25 + 0.14 + 0.09 = 0.48